2. Frieden Companys contribution format income statement for the most recent month is given below Sales (41,000 units) Variable expenses $ 861,000 602,700 Contribution margin Fixed expenses 258,300 206,640 Net operating income S 51,660 The industry in which Frieden Company operates is quite sensitive to cyclical movements in the economy. Thus, profits vary considerably from year to year according to general economic conditions The company has a large amount of unused capacity and is studying ways of improving profits Required: 1. New equipment has come on the market that would allow Frieden Company to automate a portion of its operations. Variable expenses would be reduced by $6 30 per unit. However, fixed expenses would increase to a total of $464,940 each month. Prepare two contribution format income statements, one showing present operations and one showing how operations would appear if the new equipment is purchased 2. Refer to the income statements in (1) above. For both present operations and the proposed new operations, Compute a. The degree of operating leverage b. The break-even point in dollars c. The margin of safety in both dollar and percentage terms deciding whether to purchase the new equipment? (Assume that ample funds are available to make 3. Refer again to the data in (1) above. As a manager, what factor would be paramount in your mind in the purchase) 4. Refer to the original data. Rather than purchase new equipment, the marketing manager argues that the companys marketing strategy should be changed. Instead of paying sales commissions, which are included in variable expenses, the marketing manager suggests that salespersons be paid fixed salaries and that the company invest heavily in advertising. The marketing manager claims that this new approach would increase unit sales by 50% without any change in selling price, the companys new monthly fixed expenses would be $258,300, and its net operating income would increase by 25% Compute the break-even point in dollar sales for the company under the new marketing strategyanswer all of the questions and show the solution  
Answer #1

Solution

Frieden Company

  1. Two contribution format income statements, one showing present operations and one showing how operations would appear if the new equipment is purchased:

Frieden Company

Contribution format income statements

Present Operations

Proposed Operations

Sales

$861,000

$861,000

Variable expenses

$602,700

$344,400

Contribution Margin

$258,300

$516,600

Fixed expenses

$206,640

$464,940

Net operating income

$51,660

$51,660

  1. Computation of
  1. Degree of operating leverage –

Degree of operating leverage = contribution margin/net operating income

For Present operations,

DOL = $258,300/$51,660 = 5

For Proposed Operations (Purchase of new equipment),

DOL = $516,600/51,660 = 10

  1. The break-even point in dollars –

Break-even point in dollars = fixed cost/contribution margin ratio

Contribution margin ratio = (contribution/sales) x 100

For Present operations,

Contribution margin ratio = 258,300/861,000 = 30%

Fixed cost = $206,640

Break-even point in dollars = 206,640/30% = $688,800

For Proposed Operations,

Contribution margin ratio = 516,600/861,000 = 60%

Fixed cost = 464,940

Break-even point in dollars = 464,940/60% = $774,900

  1. Margin of safety in both dollars and percentage –

Margin of safety in dollars = actual sales – BEP sales

MOS % = margin of safety in dollars/actual sales

Present operations,

MOS = 861,000 – 688,800 = $172,200

MOS % = 172,200/861,000 = 20%

For Proposed operations,

MOS = 861,000 – 774,900

= $86,100

MOS% = 86,100/861,000 = 10%

  1. Factors paramount to the decision of purchasing the new equipment:
  1. The industry’s sensitivity to the cyclical movements in the economy indicates that since the purchase of new equipment would increase the CM ratio, the high economic activity would indicate strong sales and returns. However, when   the economic activity slows down and sales fall, the investment in new equipment would push losses further down.
  2. Also, the company already has unused capacity and the purchase of new equipment would increase the unused capacity, which might cause a financial burden if sales are affected by the industry’s sensitivity to the slower economic activity.

  1. Computation of the break-even point in dollar sales under the new marketing strategy:

Increase in unit sales = 41,000 + 50% 41,000 = 61,500

Sales price per unit = $861,000/41,000 = $21

Sales in dollars = $21 x 61,500 = $1,291,500

New monthly fixed expenses = $258,300

Increase in net operating income by 25%,

= 51,660 + 25% x 51,660 = $64,575

Hence contribution margin = 258,300 + 64,575 = $322,875

Contribution margin ratio = 322,875/1,291,500 = 25%

Break-even point in dollar sales = 258,300/25% = $1,033,200